OpenStudy (mathhelp346):

\[2m+5=\left(\begin{matrix}m \over5\end{matrix}\right)-4\] Explain how you got your answer too @Mathematics

OpenStudy (asnaseer):

\[2m+5=\frac{m}{5}-4\]multiply both sides by 5 to get:\[10m-25=m-20\]\[10m-m=25-20\]\[9m=5\]\[m=\frac{5}{9}\]

OpenStudy (asnaseer):

oops - I made an error in the sign above

OpenStudy (mathhelp346):

?

OpenStudy (asnaseer):

3rd line should be: \[10m+25=m-20\]\[10m-m=-25-20\]\[9m=-45\]\[m=-5\]

OpenStudy (mathhelp346):

thanks

OpenStudy (asnaseer):

yw